I need help how do you right this in standard form ?
1)f(x)=4-x+x^2

Answer by stanbon(75887)

right this in standard form ?
f(x)=4-x+x^2
-----
f(x) = x^2-x+4 is a standard form.
-----
Another standard form:
x^2-x+(1/2)^2 = y-4+(1/2)^2
----------
(x-(1/2))^2 = y - (15/4)
Or
y = (x-(1/2))^2 + (15/4)
=============================
Cheers,
Stan H.